OpenStudy (anonymous):

How do I find the missing lengths of a right triangle if I only know the hypotenuse which is 20

OpenStudy (anonymous):

a^2 + b^2 =20^2 let b=20-a so a^2 + (20-a)^2= 400

OpenStudy (anonymous):

then u solve for a and b=20-a

OpenStudy (anonymous):

So it could be like 5 + 15 = 20 - 5?

OpenStudy (anonymous):

no just use the equation on top to solve for variable a first

OpenStudy (anonymous):

actually wait....sorry i dont think b=20-a

OpenStudy (anonymous):

Oh ok I get it thanks!

OpenStudy (anonymous):

oh ok sure